Find the polynomial p(x) with given zeros 2,-4,5
Answer by stanbon(75887)

Find the polynomial p(x) with given zeros 2,-4,5
------
p(x) = (x-2)(x+4)(x-5)
---
p(x) = (x^2+2x-8)(x-5)
---
p(x) = x^3-2x^2-5x-5x^2-10x+40
----
p(x) = x^3-7x^2-15x+40
=================
cheers,
Stan H.